OpenStudy (gabylovesyou):

For which value of x does the graph of y = 3x2 − 2x − 5 cross the x-axis?

OpenStudy (gabylovesyou):

@galacticwavesXX

OpenStudy (anonymous):

this means solve \[3x^2-2x-5=0\]

OpenStudy (anonymous):

luckily this one factors as \[(x+1)(3x-5)=0\] so it is not too bad to solve

OpenStudy (gabylovesyou):

-1 ??

OpenStudy (anonymous):

lol just like @satellite73 the equation is factor-able and has been factored for you. there should be 2 values of x

OpenStudy (anonymous):

there is \[(x+1)=0\] and \[(3x-5)=0\] solve for the x values You already got the first one, now solve for the 2nd

OpenStudy (gabylovesyou):

5/3

OpenStudy (anonymous):

:D there ya go!!

OpenStudy (gabylovesyou):

my choices i only have -1 so ill put that one :)

OpenStudy (anonymous):

ohh okay i see what you are talking about. Put the choices down next time too, to better understand the question

OpenStudy (gabylovesyou):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (gabylovesyou):

@galacticwavesXX

OpenStudy (anonymous):

Is the 1st part the 3 options?

OpenStudy (gabylovesyou):

i guess so :p

OpenStudy (anonymous):

its the 2nd one which is \[b^2-4ac=0\] now use the quadratice equation given an plug it in to the formula above you get \[7^2-4(2)(3)=0\] then you have \[25\] which is \[5^2\] so equation is factor-able

OpenStudy (anonymous):

lol my bad i thought i had posted this up

OpenStudy (gabylovesyou):

lmao!!! what is the answer? 5^2 ? or do i put 25 lol

OpenStudy (anonymous):

hmmmm...how do they want you to put down the answer?

OpenStudy (gabylovesyou):

its a fill in the blank ..For Fill in the blank items, you may type in any digits between 0 and 9, – for a negative sign, / for a fraction bar, or a decimal point. Do NOT type in commas, dollar signs, percent signs, variables, or spaces. If there are extra spaces at the end of your answer, the computer will mark that problem incorrect.

OpenStudy (gabylovesyou):

so it cant be 25 since its numbers 0 - 9

OpenStudy (gabylovesyou):

@galacticwavesXX

OpenStudy (anonymous):

lol ok then its 5...no It Has to be 5!

OpenStudy (gabylovesyou):

let me show u my question... just to be sure

OpenStudy (anonymous):

i guarantee its 5. since its 0 to 9 plus 25 is 5^2 and 5 can be factored